Two vehicles are approaching an intersection. One is a 2800 kg pick-up traveling at 15.0 m/s from) west to east, and the other is 1750 kg sedan moving north to south at 23.0 m/s. (a) What are the


magnitude of the individual momentum of each vehicle? (b) What is the magnitude and direction of the total momentum of the system? (c) What is their total kinetic energy?

1
Expert's answer
2022-04-21T13:02:39-0400

a)

"p_1=m_1v_1=42000~\frac{kg\cdot m}s,"

"p_2=m_2v_2=40250~\frac{kg\cdot m}s,"

b)

"p=\sqrt{p_1^2+p_2^2}=58173~\frac{kg\cdot m}s,"

c)

"E=\frac{p_1v_1}2+\frac{p_2v_2}2=778~kJ."